# TOTEM: TOkenized Time Series EMbeddings for General Time Series Analysis

### v1

@@ 3 Method
Table 2: Generalist Imputation ()normal-↓($\downarrow$)( ↓ ). TOTEM & GPT2 simultaneously train on all in domain datasets, 3 seeds each. A. In-Domain Performance. TOTEM has the highest AvgWins at 58.3%. B. Zero-Shot Performance. We test on unseen datasets zero-shot. TOTEM again has the highest AvgWins at 80.0%.
[t] A. In-Domain Performance Model TOTEM GPT2 Metric MSE MAE MSE MAE W 12.5% 0.029 0.060 0.029 0.045 25% 0.030 0.060 0.033 0.048 37.5% 0.032 0.062 0.037 0.054 50% 0.036 0.067 0.043 0.061 E 12.5% 0.065 0.171 0.080 0.186 25% 0.071 0.179 0.091 0.197 37.5% 0.080 0.189 0.108 0.213 50% 0.095 0.205 0.132 0.236 m1 12.5% 0.041 0.132 0.052 0.141 25% 0.044 0.135 0.065 0.154 37.5% 0.048 0.139 0.085 0.171 50% 0.058 0.152 0.117 0.196 m2 12.5% 0.040 0.125 0.029 0.095 25% 0.041 0.126 0.033 0.101 37.5% 0.043 0.129 0.038 0.110 50% 0.048 0.136 0.045 0.121 h1 12.5% 0.100 0.201 0.113 0.217 25% 0.108 0.209 0.131 0.231 37.5% 0.122 0.220 0.153 0.247 50% 0.144 0.237 0.182 0.266 h2 12.5% 0.075 0.175 0.067 0.155 25% 0.076 0.177 0.071 0.160 37.5% 0.093 0.195 0.077 0.167 50% 0.089 0.192 0.086 0.179 AvgWins 58.3% 43.8%
B. Zero-Shot Performance
Model
TOTEM
GPT2
@@ B. Zero-Shot Performance
0.082
0.193
0.128
0.234
AvgWins
80.0% 20.0%
4 Experimental Setup
Through experiments in imputation (§5), anomaly detection (§6), and forecasting (§7), our goal is to explore the efficacy of TOTEM on standard benchmark datasets and tasks, and domain general settings. To briefly refresh: specialist refers to training on a single domain (Tables 1, 3, 5). Generalist refers to training on multiple domains (Tables 2, 4, 6). Finally, in-domain refers to testing on the training domain, and zero-shot to testing on a separate domain from training.
Figure 2: Forecaster Modeling. The forecasting task requires modeling beyond the VQVAE. We leverage TOTEM’s pretrained, learnt, discrete codes as a the input data representation and train a transformer encoder. We add positional embeddings along the time dimension, and use linear layers before the final output as well as to un-normalize the resulting forecast.
For all experiments & models, we run three seeds and report the mean; standard deviations are reported in the Appendix A. Following the field standard, we bold the best metric in all tables. Evaluation metrics differ across tasks. We report mean squared error MSE ($\downarrow$), mean absolute error MAE ($\downarrow$), precision P ($\uparrow$), recall R ($\uparrow$), and F1 score ($\uparrow$); ($\downarrow$) means lower is better, ($\uparrow$) means higher performance is better. Given the varied metrics we calculate the average number of best results, or AvgWins , for each method and highlight the best, second best, and third best methods.
Baselines. We compare to two families of approaches: methods designed for multiple tasks (multitask) – TOTEM belongs in this category – and methods designed for a specific task (singletask), and are adapted to other tasks.
We compare against two recent multitask models, the transformer based GPT2 (Zhou et al., 2023) and the convolutional TimesNet [TiNet] (Wu et al., 2022a). For singletask models we compare against PatchTST [Patch] (Nie et al., 2022), ETSFormer [ETS] (Woo et al., 2022), Fedformer [FED] (Zhou et al., 2022), Non-stationary trans. [Stat] (Liu et al., 2022b), Autoformer [Auto] (Wu et al., 2021), Informer [Inf] (Zhou et al., 2021), Reformer [Re] (Kitaev et al., 2020), LightTS [LiTS] (Zhang et al., 2022), DLinear [DLin] (Zeng et al., 2023), Anomaly trans. [ATran](Xu et al., 2021), Pyraformer [Pyra] (Liu et al., 2021), LogTrans. [LogTr] (Li et al., 2019), Trans. [Trans] (Vaswani et al., 2017), Crossformer [Cross] (Zhang & Yan, 2022), TiDE (Das et al., 2023), RLinear [RLin] (Li et al., 2023), SciNet [SCi] (Liu et al., 2022a), & iTrans. [iTrans] (Liu et al., 2023).
Datasets. We leverage 12 benchmark datasets: weather [W], electricity [E], traffic [T], ETTm1 [m1], ETTm2 [m2], ETTh1 [h1], ETTh2 [h2], SMD, MSL, SMAP, SWAT, PSM that are commonly used for imputation, anomaly detection and forecasting (Zhou et al., 2023; Wu et al., 2022a; Xu et al., 2021; Zhang & Yan, 2022; Nie et al., 2022). For the zero shot settings, we leverage 5 benchmark datasets: neuro2 [N2], neuro5 [N5] (Peterson et al., 2022), and saugeen river flow [R], U.S. births [B], and sunspot [S] (Godahewa et al., 2021). 17 datasets in total.
5 Imputation
In imputation, models intake a masked time series $\mathbf{x_{m}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$, and then reconstruct and impute $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ We experiment with four canonical masking percentages at $12.5\%,25\%,37.5\%,50\%$, and report MSE and MAE .
Specialist. In Table 1 we compare TOTEM to baselines. All models are trained and evaluated on the same dataset (in-domain). TOTEM has the highest AvgWins with 52.1%, followed by GPT2 at 35.4%, and TiNet at 18.8%. TOTEM performance for m1 and h1 is lower; notably these datasets are the minute and hour resampling of the same raw data respectively. We investigate and discuss TOTEM’s success across different domains in Table 9.
Generalist. In Table 2 we compare TOTEM to GPT2 (best performing models above), when both models are trained on the aggregate of W, E, m1, m2, h1, h2. We test them on the in-domain and zero-shot test sets. TOTEM outperforms GPT2 in-domain, 58.3% vs. 43.8% , and by a much larger margin in zero-shot, 80% vs. 20%.
TOTEM’s performance across all experiments demonstrate that tokens are a performant representation for imputation.
@@ 5 Imputation
Table 3: Specialist Anomaly Detection ()normal-↑($\uparrow$)( ↑ ). TOTEM has the highest AvgWins at 33.3% followed by a five-way tie between GPT2, TiNet, ATrans, ETS, and LogTr at 13.3%. Some prior methods use the test set as a validation set for early stopping of the learning algorithm, which can inflate performance. We do not adopt this practice and train TOTEM for a set number of iterations.
@@
Table 4: Generalist Anomaly Detection ()normal-↑($\uparrow$)( ↑ ). We train TOTEM & GPT2 on all datasets and then perform in-domain and zero-shot evaluations. A. In-Domain Performance. TOTEM outperforms GPT2: 80.0% vs. 20.0%. B. Zero-Shot Performance. TOTEM again outperforms GPT2: 73.3% vs. 26.7%.
@@ B. Zero-Shot Performance
6 Anomaly Detection
In anomaly detection, models intake a corrupted time series $\mathbf{x_{corr}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and reconstruct the data $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$, where the amount of corruption is considered known, at A%. We report % Precision P ($\uparrow$), Recall R ($\uparrow$), and F1 Score ($\uparrow$).
The standard practice in machine learning, which we adopt, is to have a held out test set that is not used for tuning the model or learning algorithm. One aspect that makes comparing with several prior works challenging is that they use the test set as a validation set for early stopping of the learning algorithm, which can often inflate their performance. Despite this inconsistency, we compare our performance against these reported performances, whenever available.
Specialist. In Table 3 we evaluate TOTEM against numerous specialist baselines. TOTEM has the highest AvgWins at 33.3% followed by a five-way tie between GPT2, TiNet, ATrans, ETS, and LogTr at 13.3%.
Generalist. In Table 4 we compare generalist-trained TOTEM and GPT2. On the in-domain test sets TOTEM outperforms GPT2: 80% vs. 20%. In the zero-shot test sets TOTEM outperforms GPT2: 73.3% vs. 26.7%.
TOTEM’s AvgWins across the specialist and generalist settings demonstrate that tokens are a performant representation for anomaly detection.
Table 5: Specialist Forecasting ()normal-↓($\downarrow$)( ↓ ). TOTEM has the best AvgWins (28.6%), followed by iTrans (26.8%). Notably, TOTEM has first place finishes in 5 datasets, while iTrans’ first places are concentrated in only electricity and traffic. All models have lookback $T_{in}=96$.

### v2

@@ D. Specialist In-Domain
Table 5: Imputation Summary. In all categories TOTEM has SOTA AvgWins . In the specialist TOTEM has $52.1\%$ AvgWins ; in generalist in domain TOTEM has $58.3\%$; in generalist zero shot TOTEM has $80.0\%$.
Since we only use 3 seeds, we run a non-parametric permutation test on the generalist models in Appendix A.6 to analyze the performance of TOTEM vs. GPT2 (Table 24), and TOTEM vs. PatchTOTEM (Table 25). We find that TOTEM statistically significantly ($p\leq 0.05$) outperforms GPT2 in terms of AvgWins on all tasks for both the in-domain and zero-shot testing paradigms. Additionally, TOTEM outperforms PatchTOTEM in a statistically significant ($p\leq 0.05$) manner for in-domain and zero-shot testing.
5.1 Imputation
In imputation, models intake a masked time series $\mathbf{x_{m}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$, and then impute the signal $\mathbf{\hat{x}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ (see Figure 11). We experiment with four canonical masking percentages at $12.5\%,25\%,37.5\%,50\%$, and report the resulting MSE and MAE.
Specialists. Figure 5A and Table 5D compare TOTEM to specialist baselines. All models are trained and evaluated on the same dataset (in-domain). TOTEM has the highest AvgWins with 52.1%, followed by GPT2 at 35.4%, and TiNet at 18.8%. TOTEM’s performance on m1 and h1 is lower, but since these datasets are the minute and hour resampling of the same raw data respectively, we expect their results to be correlated.
Generalists. Figure 5A and Tables 5B&C compare TOTEM to GPT2 (the best two models in the specialist in-domain regime) in the generalist setting, when both models are trained on the aggregate of the W, E, m1, m2, h1, and h2 datasets. We evaluate them on both the in-domain and zero-shot test sets. TOTEM outperforms GPT2 in-domain, 58.3% vs. 43.8%, and by a much larger margin zero-shot, 80% vs. 20%. TOTEM’s performance across all experiments demonstrate that tokens are a performant representation for imputation. We visualize codebook examples in Figure 13, and imputation examples in Figure 14.
5.2 Anomaly Detection
Figure 5: Anomaly Detection Results. In all cases, TOTEM has SOTA AvgWins. Vs. specialists, TOTEM has $33.3\%$; vs. generalists in-domain, TOTEM has $80.0\%$; vs. generalists zero-shot, TOTEM has $73.3\%$.
In anomaly detection, models intake a corrupted time series $\mathbf{x_{corr}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and predict which times correspond to anomalies via a binary mask $\mathbf{\hat{y}}\in\{0,1\}^{T_{\text{in}}}$, where the amount of corruption is considered known, at A(see Figure 12). We report Precision P ($\uparrow$), Recall R ($\uparrow$), and F1 Score ($\uparrow$). In the main text, we compare against the flawed baselines from prior work (see Section 4.4) for ease of comparison. We compare against a subset of 15 “correct” baselines from Wu & Keogh (2021) in Table 13. In both cases, we find that TOTEM achieves SOTA results.
Specialists. Figure 5 and Table 10 test TOTEM against specialist baselines. TOTEM has the highest AvgWins at 33.3% followed by a tie between GPT2, TiNet, ATrans, ETS, and LogTr at 13.3%.
Generalists. Figure 5 and Table 11 compare generalist-trained TOTEM and GPT2. On the in-domain and zero-shot regimes, TOTEM outperforms GPT2 80% to 20% and 73.3% to 26.7% respectively. TOTEM’s AvgWins across the specialist and generalist settings demonstrate that tokens are a performant representation for anomaly detection.
5.3 Forecasting
In forecasting, models intake a time series $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and predict future readings $\mathbf{y}\in\mathbb{R}^{S\times T_{\textrm{out}}}$, where $S$ is the sensor dimension and $T_{\textrm{in}},T_{\textrm{out}}$ signify the durations of the preceding and succeeding time series, respectively. All models have a lookback window of $T_{\textrm{in}}=96$, with prediction lengths $T_{\textrm{out}}=\{96,192,336,720\}$. Results for baselines are from Liu et al. (2023). We run GPT2 with $T_{\textrm{in}}=96$ as Zhou et al. (2023) originally use inconsistent dataset-specific lookback lengths. See Figure 7 for a summary.
Specialists. Figure 7 and Table 14 show that TOTEM achieves the highest AvgWins at 28.6% followed by iTrans at 26.8%. In particular, TOTEM has first-place finishes in five datasets while iTrans’ first-place finishes are concentrated in only the electricity and traffic datasets.
Generalists. Figure 7 and Table 15 compare the generalist-trained TOTEM and GPT2 models. TOTEM outperforms GPT2 in both the in-domain (67.9% vs. 33.9%) and zero-shot (90.0% vs. 12.5%) regimes. TOTEM’s AvgWins across both regimes show that tokens are a performant representation for forecasting.
Figure 6: Forecasting Summary. In all categories TOTEM has SOTA AvgWins . In the specialist TOTEM has $28.6\%$; in generalist in domain TOTEM has $67.9\%$; in generalist zero shot TOTEM has $90.0\%$.
Figure 7: Discrete Token Ablation. In all categories, the discrete token representation (TOTEM) has SOTA AvgWins over the patch representation (PatchTOTEM).
6 Ablations
We present 4 ablation studies: (i) testing tokens vs. patches for a fixed TOTEM architecture, (ii) testing tokens vs. patches using both transformer and MLP forecasters, (iii) a codebook size study, and (iv) a study of TOTEM’s zero-shot performance when trained on datasets of different sizes.
Tokens vs. Patches. The experiments in Section 5 show that the combination of discrete tokenization and TOTEM’s generalist architecture achieve SOTA performance. We now fix the architecture while varying only the representation (TOTEM vs. PatchTOTEM) on a forecasting task to test what proportion of the performance is attributable to tokenization. We find that in all testing regimes used in the main results, TOTEM greatly outperforms PatchTOTEM, with 67.9% vs. 39.3% AvgWins in the specialist in-domain regime, 78.6% vs. 23.2% AvgWins in the generalist in-domain regime, and 67.5% vs. 35.0% AvgWins in the generalist zero-shot regime (see Figure 7 and Table 21).
Figure 8: Discrete Token vs. Patches with MLP. For both the transformer (left) and MLP (right) the discrete token representation (TOTEM) outperforms the patch respresentation (PatchTOTEM).
Downstream Architecture Study. In Figure 8 & Table 21, we explore the effect of discrete tokens vs. patches for each of two common downstream forecasting models: the transformer encoder introduced in Section 3.3 and Figure 4, and an MLP (Ekambaram et al., 2023; Das et al., 2023a; Zeng et al., 2023). The MLP has 3-layers ReLU activations, uses dropout with $p=0.1$ after the second layer, and concludes with a layernorm; this architecture is modeled after similar architectures in the literature like Das et al. (2023a). The patch-based MLP takes in an uncompressed time series. We find that for both the MLP and transformer architectures, the discrete token representation outperforms the patch representation (in the transformer $67.9\%$ to $39.3\%$ AvgWins and MLP $66.1\%$ to $37.5\%$ AvgWins). This shows that TOTEM’s strength in forecasting is not due to the strength of the transformer forecaster, but because of the choice to use discrete tokens.
Figure 9: Codebook Size Ablation. As the codebook size $K$ increases, the reconstruction loss of the VQVAE decreases on a variety of datasets.
Figure 10: Dataset Size Study. As expected, the generalist has the highest zero-shot performance at $85.0\%$ AvgWins, but the electricity specialist outperforms the traffic specialist even with a smaller training dataset. This confirms that dataset diversity may be more important than dataset scale for generalization.
Codebook Size. In Figure 10, we explore the effect of the codebook size $K$ on the VQVAE’s reconstruction performance. As expected, we find that as $K$ increases from 32 to 256 to 512, the reconstruction performance improves. However, for downstream tasks like forecasting, it is more parsimonious to model interactions between fewer codewords. Thus, we elect to use $K=256$ codewords, as the reconstruction performance is similar to that of $K=512$. We note that the the average generalist codebook error (see Table 21D), is substantially lower than the corresponding downstream forecasting error, demonstrating that a larger proportion of error is attributable to the difficulty of the forecasting task rather than poor reconstruction. This provides evidence that time series can have a single unified representation across multiple domains, akin to BPE in language modeling. We note that this same trend holds for the specialist models as well.
Dataset Size Study. One natural question is whether TOTEM’s strong generalization performance is driven by the size of the dataset or the diversity of the training samples. We study this in a minimal setting by comparing the TOTEM generalist model against two TOTEM specialists trained on the two largest domain-specific datasets: traffic (10.2M examples) and electricity (5.8M examples). As expected, the results in Figure 10 show that the TOTEM generalist significantly outperforms the two specialists in the zero-shot setting. However, the electricity specialist outperformed the traffic specialist even though the training dataset was about half the size. This provides some preliminary evidence that simply training on more data is insufficient for achieving generalization - the types of data are also crucial. For related exploratory studies on generalist models, see Appendix A.7.
7 Conclusion
We present TOTEM: a simple, performant tokenizer that is designed to learn domain-agnostic discrete representations for time series data, paving the way for time series foundation models. TOTEM demonstrates strong in-domain and zero-shot capabilities versus a large array of both generalist and specialist baselines across dozens of domains and datasets over hundreds of seeded experiments. Overall, TOTEM unlocks domain generalization while performing at or above existing SOTA levels, demonstrating the potential of adopting training and modeling techniques from language and vision modeling for time series modeling.
There are many exciting directions for future work. First, our proposed architectural design decisions were very simple, which suggests that there are many possible performant extensions. Further, while we have collated millions of existing time series, TOTEM’s promising initial results suggest that scaling up the generalist training dataset size by an order of magnitude or more could unlock true domain- and task-agnostic generalizability. Such followup works could allow a more systematic study of the relationships between generalist data representations, token length, data size, and domain diversity.
8 Broader Impact Statement
9 Acknowledgments
We thank Albert Hao Li for helpful discussions and edits, Addison Hu for insights into statistical modeling, and Angela Gao and Jack Wilding for discussions surrounding applications to earthquake data.
Appendix A Appendix
A.1 Dataset.
Dataset Sampling Rate Number Sensors
Imputation & Forecasting Training Sets
Weather Every 10 min 21
Traffic Every hour 862
Electricity Every hour 321
Etth1, ETTh2 Every hour 7
Ettm1, ETTm2 Every 15 min 7
Anomaly Detection Training Sets
SMD (Sever Machine) Every min 38
MSL (Mars Rover) Every min 55
SMAP (Soil Moisture) Every min 25
SWAT (Water Treatment) Every sec 51
PSM (Pooled Server) Every min 25
Zero Shot Testing Sets for Imputation, Forecasting & Anomaly Detection
Neuro2 Every 0.002 sec 72
Neuro5 Every 0.002 sec 106
Saugeen River Flow Every day 1
US Birth Rate Every day 1
Sunspot Every day 1
Table 6: Dataset Information Table. Notably no sampling rate or sensor number is shared between the training sets and testing sets for any task.
Task Training Domains Training Domains Zero Shot Testing
Explored (Main Paper) Explored (Appendix) Domains Explored
Imputation Weather, Electricity, Healthcare Neuroscience (ECoG), River Flow,
Transformer Temperature U.S. Birth Rate, Sunspot
Anomaly Server Machines, Mars Science Lab, Insect Feeding, Walking Acceleration, Neuroscience (ECoG), River Flow,
Detection Soil Moisture, Water Treatment Air Temperature, 3D Gait Phase, U.S. Birth Rate, Sunspot
Heart Beat (including ECG datasets),
Parkinson’s Asymmetry, Tilt Table Beat,
Internal Bleeding, Accelerometer on Whale,
NASA SpaceCraft Increase Rate
Forecasting Traffic, Weather, Electricity, Tourism, Imports & Exports, Neuroscience (ECoG), River Flow,
Transformer Temperature Real Estate, etc. [aggregated in the W4, U.S. Birth Rate, Sunspot
W3, etc. datasets], Demographics
Table 7: Domain Diversity Table. Here we list various domains that are used for each task across training and testing in both the main paper and Appendix.
A.2 Imputation.
Figure 11: Imputation Visualization. The VQVAE architecture does not change for the imputation task. The data passed in has a mask applied to it so that the VQVAE solves the task of reconstruction and imputation simultaneously.
Table 8: Means & Stds. for the Imputation Task. A. is the TOTEM specialist, B. is the TOTEM generalist, C. is the GPT2 generalist which we setup to run in a generalist manner.
A. TOTEM - Specialist Imputation ()($\downarrow$)( ↓ )
@@ B. TOTEM - Generalist Imputation ()($\downarrow$)( ↓ )
$0.082\pm 0.0010$
$0.193\pm 0.0015$
C. GPT2 - Generalist Imputation ()($\downarrow$)( ↓ )
Metric
MSE
MAE
W
12.5%
$0.029\pm 0.0000$
$0.045\pm 0.0006$
25%
$0.033\pm 0.0006$
$0.048\pm 0.0006$
37.5%
$0.037\pm 0.0006$
$0.054\pm 0.0012$
50%
$0.043\pm 0.0012$
$0.061\pm 0.0017$
E
12.5%
$0.008\pm 0.0020$
$0.186\pm 0.0035$
25%
$0.091\pm 0.0020$
$0.197\pm 0.0025$
37.5%
$0.108\pm 0.0021$
$0.213\pm 0.0026$
50%
$0.132\pm 0.0026$
$0.236\pm 0.0026$
m1
12.5%
$0.052\pm 0.0012$
$0.141\pm 0.0016$
25%
$0.065\pm 0.0021$
$0.154\pm 0.0021$
37.5%
$0.085\pm 0.0038$
$0.171\pm 0.0026$
50%
$0.117\pm 0.0052$
$0.196\pm 0.0026$
m2
12.5%
$0.029\pm 0.0000$
$0.095\pm 0.0006$
25%
$0.033\pm 0.0006$
$0.101\pm 0.0006$
37.5%
$0.038\pm 0.0006$
$0.110\pm 0.0012$
50%
$0.045\pm 0.0006$
$0.121\pm 0.0012$
h1
12.5%
$0.113\pm 0.0012$
$0.217\pm 0.0021$
25%
$0.131\pm 0.0010$
$0.231\pm 0.0015$
37.5%
$0.153\pm 0.0012$
$0.247\pm 0.0017$
50%
$0.182\pm 0.0006$
$0.266\pm 0.0012$
h2
12.5%
$0.067\pm 0.0010$
$0.155\pm 0.0015$
25%
$0.071\pm 0.0006$
$0.160\pm 0.0015$
37.5%
$0.077\pm 0.0010$
$0.167\pm 0.0015$
50%
$0.086\pm 0.0032$
$0.179\pm 0.0038$
Zero-Shot
N2
12.5%
$0.047\pm 0.0006$
$0.145\pm 0.0015$
25%
$0.064\pm 0.0017$
$0.164\pm 0.0015$
37.5%
$0.090\pm 0.0036$
$0.191\pm 0.0032$
50%
$0.131\pm 0.0051$
$0.228\pm 0.0044$
N5
12.5%
$0.021\pm 0.0006$
$0.095\pm 0.0012$
25%
$0.028\pm 0.0006$
$0.107\pm 0.0010$
37.5%
$0.039\pm 0.0015$
$0.123\pm 0.0015$
50%
$0.055\pm 0.0015$
$0.145\pm 0.0023$
R
12.5%
$0.093\pm 0.0010$
$0.119\pm 0.0015$
25%
$0.125\pm 0.0006$
$0.134\pm 0.0026$
37.5%
$0.167\pm 0.0021$
$0.154\pm 0.0042$
50%
$0.220\pm 0.0045$
$0.182\pm 0.0057$
B
12.5%
$0.392\pm 0.0064$
$0.496\pm 0.0023$
25%
$0.444\pm 0.0071$
$0.523\pm 0.0029$
37.5%
$0.498\pm 0.0080$
$0.553\pm 0.0023$
50%
$0.591\pm 0.0700$
$0.599\pm 0.0275$
s
12.5%
$0.070\pm 0.0012$
$0.173\pm 0.0017$
25%
$0.084\pm 0.0010$
$0.189\pm 0.0015$
37.5%
$0.103\pm 0.0010$
$0.209\pm 0.0021$
50%
$0.128\pm 0.0015$
$0.234\pm 0.0021$
Table 9: Imputation on PhysioNet 2012 Dataset. We report MAE where lower is better. TOTEM has the best performance in all three scenarios of percent missing.
Method
[…]
V-Rin
0.271
0.365
0.606
BRITS
0.284
0.368
0.517
RDIS
0.319
0.419
0.613
Unconditional
0.326
0.417
0.625
CSDI
0.217
0.301
0.481
TOTEM (Ours)
0.126
0.134
0.143
A.3 Anomaly Detection.
[…]
@@ C. GPT2 - Generalist Imputation ()($\downarrow$)( ↓ )
Table 10: Specialist Anomaly Detection ()($\uparrow$)( ↑ ). TOTEM has the highest AvgWins at 33.3% followed by a five-way tie between GPT2, TiNet, ATrans, ETS, and LogTr at 13.3%. Some prior methods use the test set as a validation set for early stopping of the learning algorithm, which can inflate performance. We do not adopt this practice and train TOTEM for a set number of iterations.
@@
Table 11: Generalist Anomaly Detection ()($\uparrow$)( ↑ ). We train TOTEM & GPT2 on all datasets and then perform in-domain and zero-shot evaluations. A. In-Domain Performance. TOTEM outperforms GPT2: 80.0% vs. 20.0%. B. Zero-Shot Performance. TOTEM again outperforms GPT2: 73.3% vs. 26.7%.
@@ B. Zero-Shot Performance
Table 12: Means & Stds. for the Anomaly Detection Task. A. is the TOTEM specialist, B. is the TOTEM generalist, C. is the GPT2 generalist which we setup to run in a generalist manner.
A. TOTEM - Specialist Anomaly Detection ()($\uparrow$)( ↑ )
Mean $\pm$ Std
F1
SMD
$0.7962\pm 0.0137$
MSL
$0.8258\pm 0.0052$
SMAP
$0.9402\pm 0.0008$
SWAT
$0.9427\pm 0.0006$
PSM
$0.9587\pm 0.0008$
R
SMD
$0.7606\pm 0.0207$
MSL
$0.8285\pm 0.0071$
SMAP
$0.9404\pm 0.0013$
SWAT
$0.9591\pm 0.0012$
PSM
$0.9421\pm 0.0004$
P
SMD
$0.8354\pm 0.0054$
MSL
$0.8232\pm 0.0033$
SMAP
$0.9400\pm 0.0004$
SWAT
$0.9268\pm 0.0003$
PSM
$0.9758\pm 0.0012$
B. TOTEM - Generalist Anomaly Detection ()($\uparrow$)( ↑ )
Mean $\pm$ Std
F1
SMD
$0.7864\pm 0.0386$
MSL
$0.8329\pm 0.0020$
SMAP
$0.9251\pm 0.0014$
SWAT
$0.9437\pm 0.0005$
PSM
$0.9578\pm 0.0002$
N2
$0.5129\pm 0.0397$
N5
$0.5128\pm 0.0390$
R
$0.4939\pm 0.0625$
B
$0.4915\pm 0.0229$
S
$0.5217\pm 0.0418$
R
SMD
$0.7207\pm 0.0565$
MSL
$0.8296\pm 0.0046$
SMAP
$0.9148\pm 0.0020$
SWAT
$0.9613\pm 0.0010$
PSM
$0.9390\pm 0.0004$
N2
$0.7688\pm 0.0594$
N5
$0.7684\pm 0.0582$
R
$0.7049\pm 0.0825$
B
$0.7371\pm 0.0340$
S
$0.7736\pm 0.0581$
P
SMD
$0.8666\pm 0.0114$
MSL
$0.8364\pm 0.0014$
SMAP
$0.9356\pm 0.0009$
SWAT
$0.9268\pm 0.0001$
PSM
$0.9774\pm 0.0002$
N2
$0.3849\pm 0.0299$
N5
$0.3848\pm 0.0294$
R
$0.3802\pm 0.0502$
B
$0.3686\pm 0.0172$
S
$0.3935\pm 0.0325$
C. GPT2 - Generalist Anomaly Detection ()($\uparrow$)( ↑ )
Mean $\pm$ Std
F1
SMD
$0.7973\pm 0.0326$
MSL
$0.8017\pm 0.0205$
SMAP
$0.6705\pm 0.0041$
SWAT
$0.8962\pm 0.0016$
PSM
$0.9047\pm 0.0759$
N2
$0.3902\pm 0.0596$
N5
$0.4219\pm 0.0047$
R
$0.3614\pm 0.0204$
B
$0.2081\pm 0.0462$
S
$0.3812\pm 0.0621$
R
SMD
$0.7342\pm 0.0559$
MSL
$0.7848\pm 0.0277$
SMAP
$0.5342\pm 0.0051$
SWAT
$0.8753\pm 0.0033$
PSM
$0.8776\pm 0.0624$
N2
$0.3369\pm 0.0592$
N5
$0.3677\pm 0.0498$
R
$0.2966\pm 0.0218$
B
$0.1767\pm 0.0426$
S
$0.3183\pm 0.0648$
P
SMD
$0.8744\pm 0.0029$
MSL
$0.8195\pm 0.0130$
SMAP
$0.9001\pm 0.0007$
SWAT
$0.9183\pm 0.0006$
PSM
$0.9339\pm 0.0925$
N2
$0.4643\pm 0.0561$
N5
$0.4958\pm 0.0396$
R
$0.4630\pm 0.0139$
B
$0.2533\pm 0.0498$
S
$0.4772\pm 0.5000$
Table 13: Extra Anomaly Detection ()($\uparrow$)( ↑ ). We present the the Adj. F1 metric the table (higher is better), then calculate the AvgWins . The selection criteria for the 15 datasets from (Wu & Keogh, 2021; Goswami et al., 2024) was the following. First, based only on the names in (Goswami et al., 2024), it was often ambiguous which data file was used. In these cases, we excluded the dataset. Second, we had difficulty verifying whether the default train/val/test ratios specified in the (Goswami et al., 2024) code matched what was reported. We found for the majority of datasets that the defaults resulted in test sets with no anomalies, when anomalies should be present. These were also excluded. From the results we could obtain, TOTEM matches or beats all other methods.
Model
TOTEM
ATran
MNT-0
MNT-LP
DGHL
GPT2
TiNet
CIMIS44AirTemperature3
73.8
6.0
100.0
98.0
50.0
18.0
47.0
GP711MarkerLFM5z4
96.7
76.0
69.0
97.0
31.0
48.0
90.0
InternalBleeding5
100.0
94.0
100.0
100.0
100.0
92.0
100.0
MesoplodonDensirostris
99.4
100.0
91.0
84.0
79.0
100.0
100.0
TKeepSecondMARS
100.0
83.0
95.0
100.0
16.0
12.0
95.0
WalkingAceleration5
100.0
99.0
100.0
100.0
91.0
87.0
93.0
insectEPG2
100.0
12.0
11.0
23.0
14.0
81.0
96.0
ltstdbs30791AS
100.0
100.0
100.0
100.0
100.0
100.0
100.0
park3m
67.2
15.0
56.0
64.0
20.0
63.0
93.0
s20101mML2
100.0
69.0
65.0
71.0
15.0
5.0
8.0
sddb49
99.8
89.0
100.0
100.0
88.0
94.0
100.0
sel840mECG1
99.5
16.0
61.0
66.0
28.0
21.0
36.0
sel840mECG2
86.8
15.0
36.0
39.0
32.0
28.0
21.0
tiltAPB2
68.5
92.0
96.0
98.0
36.0
83.0
38.0
tiltAPB3
23.4
17.0
48.0
85.0
3.0
5.0
9.0
AvgWins
53.5%
13.3%
33.3%
53.5%
13.3%
13.3%
33.3%
Avg. Best Adj. F1
87.7
58.9
75.2
81.7
46.9
55.8
68.4
A.4 Forecasting.
Table 14: Specialist Forecasting ()($\downarrow$)( ↓ ). TOTEM has the best AvgWins (28.6%), followed by iTrans (26.8%). Notably, TOTEM has first place finishes in 5 datasets, while iTrans’ first places are concentrated in only electricity and traffic. All models have lookback $T_{in}=96$.